solve the diffrential equation x dy/dx + y = x^2y^6


1
Expert's answer
2021-02-24T06:50:32-0500

"x\frac{dy}{dx}+y=x^2y^6"

Solution:

Divide the left and right sides of the equation by "y^6" :

"\frac{x}{y^6}\frac{dy}{dx}+\frac{1}{y^5}=x^2"

"\frac{1}{y^5}=z" , "z'=-5\frac{y'}{y^6}" , "\frac{y'}{y^6}=-\frac{z'}{5}" .

"-\frac{xz'}{5}+z=x^2"

"xz'-5z=-5x^2"

"z=uv" , "z'=u'v+v'u"

"xu'v+xv'u-5uv=-5x^2"

"xu'v+u(xv'-5v)=-5x^2"

Let's compose and solve the system:

"\begin{cases}
 xv'-5v=0 \\
 xu'v=-5x^2
\end{cases}"

From the first equation:

"\frac{dv}{v}=\frac{5dx}{x}"

"\ln|v|=5\ln{|x|}"

"v=x^5"

Substitute "v" into the second equation:

"xu'x^5=-5x^2"

"du=-5x^{-4}dx"

"u=\frac53x^{-3}+C"

"z=uv=x^5(\frac53x^{-3}+C)=\frac53x^2+Cx^5"

"\frac{1}{y^5}=\frac53x^2+Cx^5"

Answer: "\frac{1}{y^5}=\frac53x^2+Cx^5" .